OpenStudy (bigbosssaint21):

Identify the factors of 6ab + 15a - 8b - 20. (3a + 4)(2b - 5) (3a - 4)(2b + 5) (3a - 5)(2b + 4) (3a + 5)(2b - 4)

OpenStudy (anonymous):

It's the second one

OpenStudy (anonymous):

3a * 2b = 6ab 3a * 5 = 15a -4 * 2b = -8b -4 * 5 = -20